OpenStudy (anonymous):

giving medals!!!! The yogurt shop wants to mix two toppings together to create a new topping that will sell for $7 per pound. The caramel crunch costs $3 per pound, and pecan turtles cost $8 dollars per pound. How many pounds of each item must be mixed together to get 10 pounds of this new topping?

OpenStudy (anonymous):

@superhelp101

OpenStudy (superhelp101):

3c + 8p = 7 * 10 c + p = 10 3c + 8p = 70 -3c - 3p = -30 5p = 40 p = 8 lbs c = 2 lbs

OpenStudy (anonymous):

so pecan turtles is 8 pounds and caramel crunch 2